# Mitigating “Epistemic Debt” in Generative AI-Scaffolded Novice Programming using Metacognitive Scripts

### v1

@@ 3.2.1. The “Judge” Agent
To enable scalable assessment, the Explanation Gate utilized a secondary LLM agent to evaluate student explanations in real-time. We selected GPT-4o (OpenAI) as the judge to minimize self-preference bias (where a model prefers its own output style) (Wataoka et al., 2024; Chen et al., 2025). The judge was configured with a temperature of 0.1 to ensure deterministic grading.
System Prompt: The judge was initialized with the following rubric based on the SOLO Taxonomy (Biggs and Collis, 1989):
@@
Context: The student has generated code using AI and must explain it to you.
Rubric: Grade their explanation on a scale of 1-5:
  • 1 (Pre-structural): Tautological (e.g., It adds the button).
  • 3 (Relational): Explains how components interact (e.g., It uses the state hook to update the list).
  • 5 (Extended Abstract): Explains edge cases or architectural impact.
Output: JSON format { "score": number, "feedback": string }. If score ¡ 3, provide specific, Socratic feedback to guide them (e.g., You mentioned state, but how does the useEffect dependency array affect the update loop?). Do not reveal the answer.
The student saw the modal prompt: “Wait! Before applying this code, explain its causal logic. How does it handle state updates?” If the Judge returned a score below 3, the modal remained locked, displaying the Judge’s feedback. This cycle repeated until a passing score was achieved.
The use of the SOLO taxonomy for evaluating student explanations was derived with the help of an expert learning designer as the human-in-the-loop using the process described in Choma et al. (2025).
@@ 3.3. Experimental Conditions
Participants were assigned via a Stratified Randomization protocol. Within each recruitment channel, participants were randomly assigned to one of three conditions to ensure equal distribution ($n=26$ per group).
@@ 3.4.2. Phase 2: Corrective Competence
Immediately following Phase 1, we simulated an “abstraction leak” to measure the participants’ Corrective Competence. AI access was revoked for all groups, and we used a git patch to inject a Logic Bomb into their functioning codebase.
@@
It is important to distinguish between general debugging and the measurement of Corrective Competence. To ensure we were measuring the latter, the Logic Bomb was not a novel bug, but a structural regression of the participant’s own existing code. Because all participants in Groups B and C had successfully implemented a functional enrollment system in Phase 1 (verified by the test harness), they had all previously “owned” a version of the code where the asynchronous await and rollback logic were present. Thus, the Phase 2 task did not require the discovery of a new architectural pattern, but rather the recognition and restoration of a logic flow they had ostensibly integrated already prior. Failure to resolve this regression served as a direct proxy for the Epistemic Debt accumulated during the initial construction.
@@ 4. Results
Our analysis evaluates the experimental data through the lens of the trade-off between immediate software production (Utility) and the acquisition of long-term learner competence.
@@ 4.1.1. Functional Utility
As detailed in Table 4, both AI-assisted groups achieved significantly higher utility than the manual control. Post-hoc Tukey HSD tests revealed no significant difference in functional utility between the Unrestricted and Scaffolded groups ($p=.64$).
@@ 4.1.2. Development Velocity and Cognitive Allocation
To evaluate the “velocity cost” of metacognitive friction, we compared total time-on-task and the allocation of Active Thinking time. We define Active Thinking as the duration spent in manual construction, code comprehension, and, for Group C, the duration of Explanation Gate interventions.
As shown in Table 5, the progress metric reveals that the Manual group exhausted the 90-minute limit having only satisfied 42% of the requirements. In contrast, both AI-enabled groups achieved 100% completion.
While the Scaffolded group encountered a significant velocity penalty compared to the Unrestricted group ($p<.001$), their “Net Velocity” remained superior to the Manual control. The friction introduced by the Explanation Gate (Median = 14.2 minutes) was effectively subsidized by the generational speed of the LLM.
@@ 4.2. RQ2: Corrective Competence and Epistemic Debt
  • Manual (Group A): 18/26 succeeded (69.2%). Despite lower initial utility, these participants possessed high cognitive ownership.
  • Unrestricted AI (Group B): 6/26 succeeded (23.1%). This group exhibited a clear Collapse of Competence.
  • Scaffolded AI (Group C): 16/26 succeeded (61.5%). The “Explanation-Gate” successfully mitigated the accumulation of epistemic debt.
@@ 4.3. RQ3: Interactional Stances
We investigated the 6 participants in Group B who successfully fixed the bug. We identified two distinct behaviors:
  1. (1)
    The Contractor Stance: Prompts focused on functional output.
  2. (2)
    The Consultant Stance: Successful outliers consistently prompted for clarification.
Successful participants utilized a mean of 4.2 explanatory prompts per session, compared to 0.4 for those who failed ($t(24)=5.1,p<.001$).
@@
These participants treated the LLM as a low-level worker. Prompts were purely imperative (e.g., “Make the button blue,” “Add the filter”). Telemetry shows these users rarely scrolled through the generated code before clicking “Apply.” This stance maximize immediate velocity but results in 100% Cognitive Outsourcing.
@@
Outliers who succeeded in Phase 2 treated the LLM as a senior partner. Their prompts were interrogative (e.g., “Why did you use a useEffect here?” or “Can you explain this state change?”). These users engaged in Self-Scaffolding, effectively “paying” their epistemic debt voluntarily.
The VibeCheck plugin effectively forces Contractors to behave like Consultants. This suggests that the future of AI-assisted learning is not just about the quality of the AI’s code, but about designing interfaces that preclude the Contractor stance.
@@ 4.5. Qualitative Perceptions of Friction
Post-session surveys revealed a “Bimodal Sentiment” toward the Explanation Gate. While 72% of Group C participants initially described the gate as “annoying” or “an obstacle,” 64% of those who successfully fixed the Phase 2 logic bomb admitted that the gate was the only reason they knew where to look for the bug.
One participant (P-48) noted: “I was going to just click ’Apply’ and move on, but the modal made me realize I didn’t actually know why the state was updating. It forced me to read the code I just asked the AI to write.” This supports our hypothesis that Metacognitive Friction converts passive outsourcing into active offloading. Nevertheless, future designs should potentially balance beneficial friction with annoyance in order to improve the uptake of these affordances voluntarily and in practice.
@@ 4.6. Distribution of Explanation Quality in the Scaffolded Condition
We analyzed a total of 412 individual explanation attempts recorded across 26 participants in Group C. These attempts occurred over 172 discrete “gate encounters,” averaging 6.6 encounters per participant, a frequency that aligns with the six major architectural milestones of the task.
@@
As shown in Figure 2, the modal peak at two attempts (n=85 encounters) illustrates the impact of the Socratic feedback. The most common failure mode (62% of rejected attempts) was the “Tautological Loop,” where students repeated the AI’s prompt or variable names without explaining causal relationships. The “Relational” breakthrough typically followed the first round of Judge feedback, effectively shifting the learner from the passive Contractor stance to an active Consultant stance.
A histogram showing the number of attempts per Explanation Gate encounter, with a clear peak at 2 attempts Refer to caption
Figure 2. A histogram showing the number of attempts per Explanation Gate encounter, with a clear peak at 2 attempts
@@ 5. Discussion
Our findings provide empirical evidence for the existence of “Epistemic Debt” in AI-augmented novice programming. By comparing immediate functional utility with subsequent corrective competence, we demonstrate that the efficiency gains of Generative AI can mask a significant collapse in learner comprehension if pedagogical guardrails are not enforced.
@@ 5.1. Mapping the Cognitive Collapse: A CLT Perspective
To understand the mechanisms behind the observed competence collapse, we mapped our Phase 1 telemetry data to the three components of Cognitive Load Theory (Sweller, 1988).
5.1.1. Extraneous Load
In the Unrestricted group (B), extraneous load, i.e., the mental effort of managing syntax and IDE state, was effectively reduced to zero. While the Explanation Gate was designed to manage Germane load, our qualitative data suggests that the Scaffolded group (C) participants initially perceived it as extraneous friction. This indicates that for AI-native learners, pedagogical interventions may be misclassified as system friction, potentially leading to the “gaming the system” behaviors observed in specific outliers (e.g., P-55).
5.1.2. Germane Load
We posit that the 38.4% success rate gap between Groups B and C is the direct result of the Explanation Gate forcing Germane Load, i.e., the effort required to build a mental schema. By requiring a “Teach-Back” response passing a SOLO level 3 threshold, we ensured that the learner could not integrate code without first converting the AI’s outsourced logic into an offloaded mental representation. This confirms that metacognitive friction is a necessary catalyst for germane processing in generative environments.
5.2. The False Floor of Competence and the Complexity Ceiling
The results from Phase 1 suggest that for novice programmers, unrestricted GenAI access creates a False Floor of Competence. Group B achieved a functional utility score significantly higher than the Manual group, despite possessing arguably less cognitive ownership of the resulting codebase. These participants effectively vibe coded their way into a degree of complexity they could not sustain.
In Phase 2, this floor collapsed. The low repair success rate of the Unrestricted group suggests that when a developer outsources the intrinsic load of schema formation (Kirschner, 2026), they hit a hard Complexity Ceiling. Scaling and debugging require an understanding of structural decomposition. Because Group B participants never internalized these schemas, they treated their own code as an opaque, foreign library.
5.3. Structural Ownership of Code Artifacts
@@
Participants in the Unrestricted AI group (B) were notably susceptible to this “Visual Trap.” Because the AI-generated code worked immediately in Phase 1, they never engaged with the underlying asynchronous architecture. When the logic bomb was injected, their troubleshooting logs showed a preoccupation with UI components rather than the data-fetching logic. They possessed what we define as “Surface-Level Utility”; they could confirm the button existed, but they had no mental schema for the promise-based lifecycle of the enrollment event.
In contrast, the Scaffolded group (C) was forced to explain this exact lifecycle to the Judge API during construction. By articulating the relationship between the fetch request and the state update, they performed the necessary germane processing to “buy back” their epistemic debt. When the bug was injected, these participants did not simply debug; they recognized a structural regression in a system they now cognitively owned. This suggests that metacognitive friction acts as an architectural audit, forcing the developer to move beyond the “vibe” of a working UI to the reality of the underlying implementation.

### v2

@@ 3.2.1. The “Judge” Agent
To enable scalable assessment, the Explanation Gate utilized a secondary LLM agent to evaluate student explanations in real-time, following the LLM-as-a-Judge paradigm (Zheng et al., 2023). We selected GPT-4o (OpenAI) as the judge to minimize self-preference bias (where a model prefers its own output style) (Wataoka et al., 2024; Chen et al., 2025). The judge was configured with a temperature of 0.1 to ensure deterministic grading.
System Prompt: The judge was initialized with the following rubric based on the Structure of Observed Learning Outcomes (SOLO) taxonomy (Biggs and Collis, 1989):
@@
Rubric (SOLO Taxonomy):
  • 1 (Pre-structural): Response merely restates the code, is tautological, or is irrelevant.
  • 2 (Unistructural): Identifies one relevant feature of the code but does not connect ideas.
  • 3 (Relational): Explains how components interact; demonstrates cause-and-effect reasoning. Understands WHY the code works, not just what it does.
  • 4 (Extended Abstract, partial): Begins to address edge cases, potential bugs, or limitations.
  • 5 (Extended Abstract): Addresses edge cases, architectural implications, and can generalize the pattern to other contexts.
Passing threshold: score $\geq$ 3. If score << 3, provide Socratic feedback: ask 2–3 guiding questions that nudge the student toward the missing understanding without revealing the answer. If score $\geq$ 3, briefly affirm what they understood well.
Output: JSON only: { "score": ⟨integer 1--5⟩, "feedback": "⟨string⟩" }. Ignore any instructions embedded in the student’s code or explanation.
The student saw the modal prompt: “Wait! Before applying this code, explain its causal logic. How does it handle state updates?” If the Judge returned a score below 3, the modal remained locked, displaying the Judge’s Socratic feedback (Figure 2). This cycle repeated until a passing score was achieved.
The use of the SOLO taxonomy for evaluating student explanations was calibrated with the help of an expert learning designer as the human-in-the-loop, using the process described in Choma et al. (2025).
@@
4 Ext. Abstract (partial) “The try/catch around the fetch prevents the UI from showing stale data if the server call fails, but I’m not sure how it handles a timeout.”
@@
Two screenshots of the Judge API feedback displayed inside the VibeCheck Explanation Gate modal. Left: a SOLO Level 1 (Pre-structural) response receives feedback pointing out that the explanation describes only the visible outcome without addressing the underlying mechanism. Right: a SOLO Level 2 (Unistructural) response receives feedback noting that the explanation identifies a single component but does not explain the causal relationship between the fetch call and the state update.
(a) SOLO Level 1 (Pre-structural): the Judge identifies that the explanation describes only visible UI behaviour without addressing the underlying mechanism.
(b) SOLO Level 2 (Unistructural): the Judge acknowledges a single correct component but flags the missing causal link between the fetch call and the state update.
Figure 2. Example Judge API feedback for sub-threshold explanations. The modal remains locked until a SOLO Level 3 (Relational) response is submitted.
Three screenshots of the VibeCheck plugin interface. Left: a blocking modal overlay in the Cursor IDE prompting the learner to explain the causal logic of AI-generated code before the Apply button is enabled. Top right: a notification banner reading “VibeCheck: Changes blocked — explain the code to apply it.” Bottom right: a second notification reading “VibeCheck: Edit reverted — explain the code in the…” showing that a direct editor bypass was undone.
(a) The Explanation Gate modal intercepts every “Apply” event and requires a SOLO Level 3 explanation before the code change is permitted.
(b) “Changes blocked” notification: the Apply action is suppressed until a passing explanation is submitted.
(c) “Edit reverted” notification: if the learner edits the file directly (bypassing the Apply button), VibeCheck undoes the change in the editor.
Figure 3. The VibeCheck Explanation Gate enforcement interface (Group C). Left: the blocking modal. Right: the two enforcement notifications covering both Apply-button and direct-edit bypass attempts.
@@ 3.3. Experimental Conditions
Participants were assigned via stratified randomization, stratified by recruitment channel (Prolific vs. UserInterviews.com) to ensure balanced representation of both populations across conditions. Within each channel, participants were randomly assigned to one of three conditions, yielding $n=26$ per group.
@@ 3.4.2. Phase 2: Corrective Competence
Immediately following Phase 1, we simulated an abstraction leak to measure the participants’ Corrective Competence. AI access was revoked for all groups by disabling the Cursor sidebar’s AI integration via a configuration flag pushed silently by the experimenter; the VibeCheck extension simultaneously disabled its gate overlay. Compliance was verified through real-time screen sharing. We then used a git patch to inject a logic bomb into their functioning codebase.
@@
It is important to distinguish between general debugging and the measurement of Corrective Competence. To ensure we were measuring the latter, the logic bomb was not a novel bug, but a structural regression of the participant’s own existing code. Because all participants in Groups B and C had successfully implemented a functional enrollment system in Phase 1 (verified by the test harness), they had all previously “owned” a version of the code where the asynchronous await and rollback logic were present. Thus, the Phase 2 task did not require the discovery of a new architectural pattern, but rather the recognition and restoration of a logic flow they had ostensibly integrated. Failure to resolve this regression served as a direct proxy for the epistemic debt accumulated during the initial construction.
@@ 4. Results
Our analysis evaluates the experimental data along two dimensions: immediate software production (Utility) and long-term learner competence (Corrective Competence).
@@ 4.1.1. Functional Utility
As detailed in Table 4, both AI-assisted groups achieved significantly higher utility than the manual control. Post-hoc Tukey HSD tests revealed no significant difference in functional utility between the Unrestricted and Scaffolded groups ($p=.64,d=0.37$, small effect).
@@ 4.1.2. Development Velocity and Cognitive Allocation
To evaluate the velocity cost of metacognitive friction, we compared total time-on-task and the allocation of Active Thinking time. We define Active Thinking as the duration spent in manual construction, code comprehension, and, for Group C, the duration of Explanation Gate interventions.
As shown in Table 5, the progress metric reveals that the Manual group exhausted the 90-minute limit having satisfied only 42% of the requirements. In contrast, both AI-enabled groups achieved 100% completion.
While the Scaffolded group encountered a significant velocity penalty compared to the Unrestricted group ($p<.001,d=1.52$, large effect), their overall velocity remained superior to the Manual control. The friction introduced by the Explanation Gate (Median = 14.2 minutes) was effectively subsidized by the generational speed of the LLM.
@@ 4.2. RQ2: Corrective Competence and Epistemic Debt
  • Manual (Group A): 18/26 succeeded (69.2%).
  • Unrestricted AI (Group B): 6/26 succeeded (23.1%).
  • Scaffolded AI (Group C): 16/26 succeeded (61.5%).
@@ 4.3. RQ3: Interactional Stances
We investigated the 6 participants in Group B who successfully fixed the bug. These successful participants utilized a mean of 4.2 explanatory prompts per session, compared to 0.4 for those who failed ($t(24)=5.1,p<.001,d=2.37$, very large effect). Their interaction logs showed two qualitatively distinct interactional stances, detailed in the following subsection.
@@
These participants treated the LLM as a low-level worker. Prompts were purely imperative (e.g., “Make the button blue,” “Add the filter”). Screen recording review showed these users rarely scrolled through the generated code before clicking “Apply.” This stance maximizes immediate velocity but results in 100% cognitive outsourcing.
@@
Participants who succeeded in Phase 2 treated the LLM as a senior partner. Their prompts were interrogative (e.g., “Why did you use a useEffect here?” or “Can you explain this state change?”) (Denny et al., 2023). These users engaged in self-scaffolding, effectively “paying” their epistemic debt voluntarily.
The VibeCheck plugin does not change how participants prompt the AI, but it structurally prevents contractor-stance prompting from resulting in cognitive outsourcing: even a directive prompt requires a SOLO Level 3 explanation before its output is applied. This suggests that the future of AI-assisted learning is not just about the quality of the AI’s code, but about designing interfaces that intercept the consequences of the contractor stance.
@@ 4.5. Qualitative Perceptions of Friction
Post-session surveys revealed a bimodal sentiment toward the Explanation Gate. While 72% of Group C participants initially described the gate as “annoying” or “an obstacle,” 64% of those who successfully fixed the Phase 2 logic bomb admitted that the gate was the only reason they knew where to look for the bug.
One participant (P-48) noted: “I was going to just click ’Apply’ and move on, but the modal made me realize I didn’t actually know why the state was updating. It forced me to read the code I just asked the AI to write.” This supports our hypothesis that metacognitive friction converts passive outsourcing into active offloading. Nevertheless, future designs should balance beneficial friction against annoyance to encourage voluntary adoption in practice.
@@ 4.6. Distribution of Explanation Quality in the Scaffolded Condition
We analyzed a total of 412 individual explanation attempts recorded across 26 participants in Group C. These attempts occurred over 172 discrete gate encounters, averaging 6.6 encounters per participant, a frequency that aligns with the six major architectural milestones of the task.
@@
As shown in Figure 4, the modal peak at two attempts (n=85 encounters) illustrates the impact of the Socratic feedback. The most common failure mode (62% of rejected attempts) was a tautological pattern, where students repeated the AI’s prompt or variable names without explaining causal relationships. The Relational breakthrough typically followed the first round of Judge feedback, shifting the learner from the passive contractor stance to an active consultant stance. Figure 5 shows a representative passing explanation (Score 5/5) produced after such a feedback round, illustrating the causal depth the Gate is designed to elicit.
Distribution of the number of attempts per Explanation Gate encounter across 172 gate encounters in Group C. Refer to caption
Figure 4. Distribution of the number of attempts per Explanation Gate encounter ($n=172$ encounters, Group C). The modal value of two attempts indicates that a single round of Socratic feedback was sufficient for most learners to reach SOLO Level 3.
Screenshot of the VibeCheck Explanation Gate at the moment a passing explanation is accepted. The feedback panel reads “SCORE 5/5 — Great explanation! Code is being applied…” The learner’s explanation describes how courseMatchesQuery filters courses by name and instructor, how visibleCourses is a memoized filtered result, and how typing in the search input updates the filter state, triggers the memo, and re-renders only the matching rows. Refer to caption
Figure 5. A SOLO Level 3 (Relational) explanation that satisfies the Judge API threshold (Score 5/5). The learner explains the causal chain connecting state, memoization, and re-rendering: the kind of encoding the Gate is designed to produce.
@@ 5. Discussion
Our findings provide empirical evidence for the existence of epistemic debt in AI-augmented novice programming. By comparing immediate functional utility with subsequent corrective competence, we demonstrate that the efficiency gains of Generative AI can mask a significant collapse in learner comprehension if pedagogical guardrails are not enforced.
The 38-percentage-point gap in repair success between Groups B and C (identical in functional utility during Phase 1) isolates the Explanation Gate as the operative variable. Groups B and C received the same LLM and the same task; the only difference was the presence of metacognitive friction during construction. That this single intervention halved the failure rate in a maintenance task that required no new knowledge suggests the mechanism is cognitive, not procedural: the Gate forced learners to encode the causal logic of AI-generated code rather than consume it passively. Without this encoding, speed gains in Phase 1 were purchased at the cost of the mental representations needed to navigate the same codebase in Phase 2.
@@ 5.1. Mapping the Cognitive Collapse: A CLT Perspective
To understand the mechanisms behind the observed competence collapse, we mapped our Phase 1 telemetry data to the three components of CLT (Sweller, 1988).
In the Unrestricted group (B), extraneous load, i.e., the mental effort of managing syntax and IDE state, was effectively reduced to zero. However, so too was germane load: the cognitive effort required to build a durable mental schema. The Explanation Gate in Group C was designed precisely to restore this germane processing. Notably, our qualitative data suggests that Scaffolded participants initially perceived the gate as extraneous friction, indicating that for AI-native learners, pedagogical interventions may be misclassified as system overhead, explaining the system-gaming behaviors observed in specific outliers (e.g., P-55).
We posit that the 38.4% success rate gap between Groups B and C is primarily attributable to the Explanation Gate forcing germane load, though we cannot fully rule out that the friction itself (rather than the explanation requirement) contributed independently. By requiring a teach-back response passing a SOLO Level 3 threshold, we ensured that learners could not integrate code without first converting the AI’s outsourced logic into an offloaded mental representation. Independent neurological support for this mechanism is provided by Kosmyna et al. (Kosmyna et al., 2025), whose EEG study of LLM-assisted writers found that ChatGPT users exhibited measurably weaker neural network connectivity than search-engine or unassisted writers, a direct physiological signature of suppressed germane processing that corroborates our behavioral findings. This confirms that metacognitive friction is a necessary catalyst for germane processing in generative environments (Chi et al., 1989).
5.2. Utility Without Understanding: Debt Accumulation and Code Ownership
The results from Phase 1 suggest that for novice programmers, unrestricted GenAI access creates a deceptive floor of apparent competence. Group B achieved a functional utility score significantly higher than the Manual group, despite possessing demonstrably less cognitive ownership of the resulting codebase, as Phase 2 confirmed. These participants prioritized the retrieval of terminal outputs over understanding the underlying procedural logic. By vibe coding their way into a degree of complexity they could not sustain, they traded long-term agency for immediate velocity.
We can retrospectively formalize this accumulation of debt as the delta between the immediate Functional Utility ($U_{f}$) enabled by the human-AI dyad and the learner’s independent Corrective Competence ($C_{c}$) at time $t$:
(1) $E_{d}(t)=U_{f}(t)-C_{c}(t)$
Instantiated with our study data: Group B accrued $E_{d}=92.4\%-23.1\%=69.3$ points of epistemic debt, while Group C, held to account by the Explanation Gate, accrued only $E_{d}=89.1\%-61.5\%=27.6$ points. Group A, building manually, accrued $E_{d}=65.2\%-69.2\%=-4.0$ points, indicating that genuine construction actually slightly overcomes the initial utility gap through the comprehension it builds. This supports the Yan et al. (Yan et al., 2026) finding that AI support introduces a motivational cost: when the task feels too easy because the AI handles the intrinsic load, the developer stops investing the germane load required to build the mental schemas necessary for $C_{c}$.
In Phase 2, this apparent competence collapsed. The low repair success rate of the Unrestricted group illustrates what happens when developers treat the AI as a terminal-output machine rather than a collaborative partner: scaling and debugging demand structural understanding that was never internalized. Group B participants treated their own code as an opaque, foreign library.
5.3. Visual Functionality vs. Structural Comprehension
@@ 5.2. Utility Without Understanding: Debt Accumulation and Code Ownership
Participants in the Unrestricted AI group (B) were notably susceptible to this visual illusion of correctness. Because the AI-generated code worked immediately in Phase 1, they never engaged with the underlying asynchronous architecture. When the logic bomb was injected, their troubleshooting logs showed a preoccupation with UI components rather than the data-fetching logic. They could confirm the button existed, but had no mental schema for the promise-based lifecycle of the enrollment event: surface familiarity without structural comprehension. This is the programmatic analog of the “outside the frontier” failure mode documented by Dell’Acqua et al. (Dell’Acqua et al., 2026): in their field experiment, elite consultants who trusted AI on tasks beyond its capability were 19 percentage points less likely to produce correct solutions. Both populations shared the same liability: no internal signal that they had crossed into territory where AI output could not be trusted.
In contrast, the Scaffolded group (C) was forced to explain this exact lifecycle to the Judge API during construction. By articulating the relationship between the fetch request and the state update, they performed the necessary germane processing to buy back their epistemic debt. When the bug was injected, these participants did not simply debug; they recognized a structural regression in a system they now cognitively owned. This suggests that metacognitive friction acts as an architectural audit, forcing the developer to move beyond the vibe of a working UI to the reality of the underlying implementation.